Candy worth $.95/kg was mixed with candy worth $1.85/kg to produce a mixture worth $1.45/kg. How many kilograms of each kind of candy were used to make 27 kg of the mixture?
Answer by mananth(16946)

---------price ---------------- quantity
Candy I 0.95 ---------------- x pounds
Candy II 1.85 ------ 27 - x pounds
Mixture 1.45% ---------------- 27

0.95 x + 1.85 ( 27 - x ) = 1.45 * 27

0.95x+49.95-1.85x=39.15
0.95x-1.85x=39.15-49.95
-0.9x= -10.8
/ -0.9
x= 12 pounds $0.95 Candy I
15 pounds $1.85 Candy II